Grade 10General Physics

when going east at 10 kmph a train moving with const. velocity appears to be moving exactly north east . when my velocity is increased to 30 kmph east it appears to be moving north. a.what is the velocity of train in noth? b.what is the velocity of train along east? ​c.what speed should i move north so that train appears to move exactly south-east? please provide complete solution in details.. :)

Akshit Bansal
9 Years agoGrade 10

1 Answer

Vikas TU
9 Years ago
 We Should apply the idea of relative speed here. 
 
Vman = 10 i 
 
Vtrain = xi+yj 
 
Vtrain,rel = (x-10)i+yj 
 
x-10 = y both part ought to be equivalent to be in north east 
 
Vtrain,rel = (x-30)i+yj 
 
x-30 =0 
 
x = 30 y =20 
 
Vtrain = 30i +20j